Some help Please

given f(x)=3x-7, find f^-1(x)
thank you for your time

Answer by funmath(2873)

given f(x)=3x-7, find f^-1(x)
Let f(x)=y
y=3x-7
Swap the x and y and solve for y
x=3y-7
x+7=3y-7+7
x+7=3y
(x+7)/3=3y/3
(1/3)x+7/3=y
Let y=f^-1(x)
f^(-1)(x)=(1/3)x+7/3
Happy Calculating!!!